# Which Backbone to Use: A Resource-efficient Domain Specific Comparison for Computer Vision

### v1

@@ Abstract
In contemporary computer vision applications, particularly image classification, architectural backbones pre-trained on large datasets like ImageNet are commonly employed as feature extractors. Despite the widespread use of these pre-trained convolutional neural networks (CNNs), there remains a gap in understanding the performance of various resource-efficient backbones across diverse domains and dataset sizes. Our study systematically evaluates multiple lightweight, pre-trained CNN backbones under consistent training settings across a variety of datasets, including natural images, medical images, galaxy images, and remote sensing images. This comprehensive analysis aims to aid machine learning practitioners in selecting the most suitable backbone for their specific problem, especially in scenarios involving small datasets where fine-tuning a pre-trained network is crucial. Even though attention-based architectures are gaining popularity, we observed that they tend to perform poorly under low data finetuning tasks compared to CNNs. We also observed that some CNN architectures such as ConvNeXt, RegNet and EfficientNet performs well compared to others on a diverse set of domains consistently. Our findings provide actionable insights into the performance trade-offs and effectiveness of different backbones, facilitating informed decision-making in model selection for a broad spectrum of computer vision domains. Our code is available here: https://github.com/pranavphoenix/Backbones
@@
Good RegNetY-3.2GF RegNetY-3.2GF ConvNeXt-Tiny ShuffleNetV2 $2.0\times$ DenseNet-161 RegNetY-3.2GF EfficientNetV2-S
@@ 2 Related Works
Most of the research on architectures uses ImageNet benchmark to compare image classification performance of backbones. Pytorch Image Models [3] library benchmarks ImageNet classification performance across different backbone architectures. Visual Task Adaptation Benchmark [4] evaluated the performance of computer vision models on 19 tasks that span different domains. Brigato et al. [5] designed a benchmark for data-efficient image classification spaning various domains and showed that tuning hyper-parameters like learning rate, weight decay, and batch size results in a competitive baseline outperforming most specialized methods. Taher et al. [6] benchmarked various self-supervised pre-training methods, 14 pre-trained ImageNet models on 7 diverse medical tasks. Battle of backbones [7] benchmarks few pre-trained ImageNet backbones, including those pretrained by self-supervised learning and Stable Diffusion, across a diverse set of computer vision tasks ranging from image classification, Out-of-distribution generalisation, image retrieval and object detection. There is a lack of bench-marking on transfer learning performance of popular lightweight backbones across multiple domain-specific datasets.
@@ 3 Backbone Architectures
Since most of the practitioners use out of the shelf ImageNet pretrained models from the popular Torchvision library, we selected models from this library for our experiments. We used certain criteria to filter models keeping in mind our resource-efficiency constraint. So, we decided to choose models with number of parameters less than 30 million, which amounts to a storage space of around 100 MB. WaveMix [8] model was added to our list of backbones becasue WaveMix was shown to provide state-of-the-art (SOTA) performance in multiple image classification datasets belonging to many diverse domains such as EMNIST [9] and galaxy morpholohy datasets.
@@
ResNet [12]: ResNets are the most popular and successful backbone architectures currently in use today since its arrival almost a dacade ago. ResNet uses residual connections to allow for the training of very deep convolutional neural networks (CNN) by mitigating the vanishing gradient problem. We use ResNet-50 for our experiments.
WaveMix [8]: WaveMix is a token-mixing architecture that uses 2-dimentional discrete wavelet transform for spacial token-mixing and has been shown to provide SOTA performance in multiple image classification datasets. We use WaveMix-192/16 (level 3) for our experiments.
ConvNeXt [13]: ConvNeXt is a recent CNN architecture designed to improve upon traditional CNNs by incorporating elements from transformer models, resulting in enhanced performance and scalability. It uses depthwise convolutions, inverted bittleneck blocks and large kernels. We use ConvNeXt-Tiny in our experiments.
Swin Transformer [10]: Swin transformer was an improvement over conventional ViT which overcame the massive data requirements for training. It incorporated efficiency using hierachial representations, limiting the attention window and merging them stage by stage. We use Swin-Tiny and SwinV2-Tiny [11] for our experiments.
@@
MobileNet [16]: MobileNetV3 is a CNN architecture designed for on-device and resource-constrained environments, which combines lightweight depthwise separable convolutions with squeeze-and-excitation modules. It is a highly efficient model with improved accuracy and reduced computational complexity. We use MobileNetV3-Large in our experiments.
@@
ShuffleNet [19]: ShuffleNet is a lightweight CNN architecture designed for efficient computation on mobile devices. It uses channel shuffling and pointwise group convolution to optimize speed and accuracy. We use ShuffleNetV2 $2.0\times$ for our experiments.
@@ 4 Datasets
Since we wanted to check the performance of various backbones on data efficient fine-tuning, we decided to perform our experiments on datasets with atmost 100,000 training images. We choose 18 piblically available datasets belonging to 6 different domains such as natural, textures, remote sensing, plants, astronomy, and medical images with number of training images ranging from 1000 to 100,000 and number of classes ranging from 2 to 200. Details of datasets in each domain is given in table 3.
@@
BreakHis $40\times$ [33] Medical Images (Histopathology) 1,398 606 2
BreakHis $100\times$ [33] Medical Images (Histopathology) 1,458 632 2
BreakHis $200\times$ [33] Medical Images (Histopathology) 1,411 611 2
BreakHis $400\times$ [33] Medical Images (Histopathology) 1,276 553 2
@@ 4.3 Remote Sensing
EuroSAT [29]: The EuroSAT dataset is a benchmark dataset for land use and land cover classification, consisting of 27,000 RGB and multispectral images covering 10 classes, with images derived from Sentinel-2 satellite data. It is widely used for evaluating the performance of machine learning algorithms in remote sensing and geospatial analysis tasks.
@@ 4.6 Medical
BreakHis [33]: Breast Cancer Histopathological Database is a dataset specifically designed for the classification of breast cancer histopathological images. It contains 7,909 microscopic images of breast tumor tissue, divided into benign and malignant categories. It provides microscopic images of breast tumor tissue at four different magnification levels: $40\times$, $100\times$, $200\times$ and $400\times$. Each magnification level offers a different level of detail, allowing for a comprehensive analysis of histopathological features. The dataset is widely used for evaluating and benchmarking algorithms in medical image analysis and computer-aided diagnosis.
@@ 5 Experimental Details
Due to limited computational resources, we use early stopping where training is stopped if accuracy stops improving after 10 epochs. All experiments were done on a single 80 GB Nvidia A100 GPU. Training was done using DualOpt [35] where we used AdamW optimizer ($\alpha=0.001,\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) with a weight decay of 0.01 during the initial phase and then used SGD with a learning rate of $0.001$ and momentum =0.9=0.9 during the final phase.
Cross-entropy loss was used for image classification. Batch-size was choosen so that the entire GPU will be utilized while training. We used automatic mixed precision in PyTorch during training. Top-1 accuracy on the test set of best of three runs with random initialization is reported as a generalization metric based on prevailing protocols [36].
@@ 6 Results
ShuffleNetV2 $2.0\times$ 78.11 91.54 74.73 88.46 76.43 96.30 83.32
Table 4: Top-1 classification accuracy for fine-tuning pretrained backbones on natural image datasets. The top 3 results for each dataset are highlighted with best in green and bold, second best in lighter green, and third best in lightest green.
@@
ShuffleNetV2 $2.0\times$ 68.64 97.86 98.65 99.70 81.76 83.69
Table 5: Top-1 classification accuracy for fine-tuning pretrained backbones on different domain datasets. The top 3 results are highlighted with best in green and bold, light green, and third best in lightest green.
@@
Backbones $40\times$ $100\times$ $200\times$ $400\times$ Average
@@
ShuffleNetV2 $2.0\times$ 99.38 98.59 99.22 98.02 98.80
Table 6: Top-1 classification accuracy for fine-tuning pretrained backbones on medical datasets. The top 3 results are highlighted with best in green and bold, light green, and third best in lightest green.
@@ 7 Observations
$\bigtriangleup$ WavMix performs well in datasets where multi-resolution token-mixing aid in learning WaveMix outperforms all other models (9% increase from the second best, ConvNeXt) in galaxy morphology classification. WaveMix also performs well in medical domain performing better than all other datasets, and also maintaining the performance across different magnification. WaveMix, which uses 2D-DWT might possess inductive bias that can analyse the domains of astronomy and medical images better than other convolutional models due its multi-resolution token-mixing. We recomment using WaveMx in domains where features across different resolutions are needed for better performance. WaveMix also performs better than ConvNeXt in CIFAR-10 and CIFAR-100 datasets, which were actually low resolution images ($32\times 32$) which were resized to ($256\times 256$) for our training. WaveMix is also state-of-the-art in many low-resolution datasets such as EMNIST ($28\times 28$). So, we recommend using WaveMix for low-resolution image datasets.
@@
$\bigtriangleup$ ConvNeXt and Swin transformers perform poorly in medical domain Our results show that both the architectures perform perform really bad compared to all the other models when used in medical domain. Therefore, we suggest practioners working in medical domain to not use these models despite their high performance in other domains.

### v3

@@ Abstract
In contemporary computer vision applications, particularly image classification, architectural backbones pre-trained on large datasets like ImageNet are commonly employed as feature extractors. Despite the widespread use of these pre-trained convolutional neural networks (CNNs), there remains a gap in understanding the performance of various resource-efficient backbones across diverse domains and dataset sizes. Our study systematically evaluates multiple lightweight, pre-trained CNN backbones under consistent training settings across a variety of datasets, including natural images, medical images, galaxy images, and remote sensing images. This comprehensive analysis aims to aid machine learning practitioners in selecting the most suitable backbone for their specific problem, especially in scenarios involving small datasets where fine-tuning a pre-trained network is crucial. Even though attention-based architectures are gaining popularity, we observed that they tend to perform poorly under low data fine-tuning tasks compared to CNNs. We also observed that some CNN architectures such as ConvNeXt, RegNet and EfficientNet performs well compared to others on a diverse set of domains consistently. Our findings provide actionable insights into the performance trade-offs and effectiveness of different backbones, facilitating informed decision-making in model selection for a broad spectrum of computer vision domains. Our code is available here: https://github.com/pranavphoenix/Backbones
@@
Good RegNetY-3.2GF RegNetY-3.2GF ConvNeXt-Tiny ShuffleNetV2 2.0×2.0$\times$2.0 × DenseNet-161 RegNetY-3.2GF EfficientNetV2-S
@@ 2 Related Works
Most of the research on architectures uses ImageNet benchmark to compare image classification performance of backbones. Pytorch Image Models [3] library benchmarks ImageNet classification performance across different backbone architectures. Visual Task Adaptation Benchmark [4] evaluated the performance of computer vision models on 19 tasks that span different domains. Brigato et al. [5] designed a benchmark for data-efficient image classification spanning various domains and showed that tuning hyper-parameters like learning rate, weight decay, and batch size results in a competitive baseline outperforming most specialized methods. Taher et al. [6] benchmarked various self-supervised pre-training methods, 14 pre-trained ImageNet models on 7 diverse medical tasks. Battle of backbones [7] benchmarks few pre-trained ImageNet backbones, including those pre-trained by self-supervised learning and Stable Diffusion, across a diverse set of computer vision tasks ranging from image classification, Out-of-distribution generalisation, image retrieval and object detection. There is a lack of bench-marking on transfer learning performance of popular lightweight backbones across multiple domain-specific datasets.
@@ 3 Backbone Architectures
Since most of the practitioners use out of the shelf ImageNet pre-trained models from the popular Torchvision library, we selected models from this library for our experiments. We used certain criteria to filter models keeping in mind our resource-efficiency constraint. So, we decided to choose models with number of parameters less than 30 million, which amounts to a storage space of around 100 MB. WaveMix [8] model was added to our list of backbones because WaveMix was shown to provide state-of-the-art (SOTA) performance in multiple image classification datasets belonging to many diverse domains such as EMNIST [9] and galaxy morphology datasets.
@@
ResNet [12]: ResNets are the most popular and successful backbone architectures currently in use today since its arrival almost a decade ago. ResNet uses residual connections to allow for the training of very deep convolutional neural networks (CNN) by mitigating the vanishing gradient problem. We use ResNet-50 for our experiments.
WaveMix [8]: WaveMix is a token-mixing architecture that uses 2-dimensional discrete wavelet transform for spacial token-mixing and has been shown to provide SOTA performance in multiple image classification datasets. We use WaveMix-192/16 (level 3) for our experiments.
ConvNeXt [13]: ConvNeXt is a recent CNN architecture designed to improve upon traditional CNNs by incorporating elements from transformer models, resulting in enhanced performance and scalability. It uses depth-wise convolutions, inverted bottleneck blocks and large kernels. We use ConvNeXt-Tiny in our experiments.
Swin Transformer [10]: Swin transformer was an improvement over conventional ViT which overcame the massive data requirements for training. It incorporated efficiency using hierarchical representations, limiting the attention window and merging them stage by stage. We use Swin-Tiny and SwinV2-Tiny [11] for our experiments.
@@
MobileNet [16]: MobileNetV3 is a CNN architecture designed for on-device and resource-constrained environments, which combines lightweight depth-wise separable convolutions with squeeze-and-excitation modules. It is a highly efficient model with improved accuracy and reduced computational complexity. We use MobileNetV3-Large in our experiments.
@@
ShuffleNet [19]: ShuffleNet is a lightweight CNN architecture designed for efficient computation on mobile devices. It uses channel shuffling and point-wise group convolution to optimize speed and accuracy. We use ShuffleNetV2 2.0×2.0$\times$2.0 × for our experiments.
@@ 4 Datasets
Since we wanted to check the performance of various backbones on data efficient fine-tuning, we decided to perform our experiments on datasets with utmost 100,000 training images. We choose 20 publicly available datasets belonging to 7 different domains such as natural, textures, remote sensing, plants, astronomy, surface defect and medical images with number of training images ranging from 1000 to 100,000 and number of classes ranging from 2 to 200. Details of datasets in each domain is given in table 3.
@@
BreakHis 40×40$\times$40 × [34] Medical Images (Histopathology) 1,398 606 2
BreakHis 100×100$\times$100 × [34] Medical Images (Histopathology) 1,458 632 2
BreakHis 200×200$\times$200 × [34] Medical Images (Histopathology) 1,411 611 2
BreakHis 400×400$\times$400 × [34] Medical Images (Histopathology) 1,276 553 2
RSNA Pneumonia Detection [35] Medical Images (Radiology) 24,181 6046 2
@@ 4.2 Texture
4.3 Surface Defect
NEU Surface Defect [28]: The Northeastern University (NEU) surface defect database is a benchmark dataset for surface defect detection and classification, featuring images of six types of surface defects. The dataset includes 1,800 images with a resolution of $200\times 200$ pixels, divided into 1,200 training images and 600 testing images. Each defect type has an equal number of images, making it a well-balanced dataset for training and evaluating machine learning models.
@@ 4.4 Remote Sensing
EuroSAT [30]: The EuroSAT dataset is a benchmark dataset for land use and land cover classification, consisting of 27,000 RGB and multi-spectral images covering 10 classes, with images derived from Sentinel-2 satellite data. It is widely used for evaluating the performance of machine learning algorithms in remote sensing and geospatial analysis tasks.
@@ 4.7 Medical
BreakHis [34]: Breast Cancer Histopathology Database is a dataset specifically designed for the classification of breast cancer histopathological images. It contains 7,909 microscopic images of breast tumor tissue, divided into benign and malignant categories. It provides microscopic images of breast tumor tissue at four different magnification levels: 40×40$\times$40 ×, 100×100$\times$100 ×, 200×200$\times$200 × and 400×400$\times$400 ×. Each magnification level offers a different level of detail, allowing for a comprehensive analysis of histopathological features. The dataset is widely used for evaluating and bench-marking algorithms in medical image analysis and computer-aided diagnosis.
RSNA Pneumonia Detection Challenge Dataset [35]: The RSNA Pneumonia Detection Challenge dataset, provided by the Radiological Society of North America, consists of over 30,000 chest X-ray images annotated for the presence of pneumonia. It is designed to facilitate the development of machine learning models for pneumonia detection, promoting advancements in medical imaging analysis.
@@ 5 Experimental Details
Due to limited computational resources, we use early stopping where training is stopped if accuracy stops improving after 10 epochs. All experiments were done on a single 80 GB Nvidia A100 GPU. Training was done using DualOpt [37] where we used AdamW optimizer ($\alpha=0.001,\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) with a weight decay of 0.01 during the initial phase and then used SGD with a learning rate of $0.001$ and momentum $=0.9$ during the final phase.
Cross-entropy loss was used for image classification. Batch-size was chosen so that the entire GPU will be utilized while training. We used automatic mixed precision in PyTorch during training. Top-1 accuracy on the test set of best of three runs with random initialization is reported as a generalization metric based on prevailing protocols [38].
@@ 6.1 Fine-tuning Performance
6.2 Performance with less training data
We observe from Table 7 that the models which are performing really well at 100% of the training data are also performing better than other models at 10% and 1% of the training data. We can see from CIFAR-10 and Tiny Imaginary Dataset results that ConvNeXt, WaveMix, EfficientNet, and even RegNet are actually performing better than others even when the training data is significantly reduced. This shows that the fine-tuning superiority of these models is actually not dependent much on the amount of training data. Even with less data, they can generalize better and create better representations than other models.
Figure 1 and Figure 2 shows the performance of top 3 models with training data. We see that the difference in accuracy amoung the three models is large and significant when the training data is less (1%) and this difference reduces as the training data increases.
We observe similar behaviour for other domains in Table 8. In plant datasets, the top performing models at 100% of training data like ConvNeXt, EfficientNet, DenseNet, and RegNet are again performing better than the other models even at 10% of the training data. Similar behaviour is also observed in radiology images, where ResNeXt and ResNeXt are performing better.
@@ 6.2 Performance with less training data
ShuffleNetV2 2.0×2.0$\times$2.0 × 78.11 91.54 74.73 88.46 76.43 96.30 83.32
Table 4: Top-1 classification accuracy for fine-tuning pre-trained backbones on natural image datasets. The top 3 results for each dataset are highlighted with best in green and bold, second best in lighter green, and third best in lightest green.
@@
ShuffleNetV2 2.0×2.0$\times$2.0 × 68.64 97.86 98.65 99.70 81.76 83.69
Table 5: Top-1 classification accuracy for fine-tuning pre-trained backbones on different domain datasets. The top 3 results are highlighted with best in green and bold, light green, and third best in lightest green.
@@
Backbones 40×40$\times$40 × 100×100$\times$100 × 200×200$\times$200 × 400×400$\times$400 × Average
@@
ShuffleNetV2 2.0×2.0$\times$2.0 × 99.38 98.59 99.22 98.02 98.80
Table 6: Top-1 classification accuracy for fine-tuning pre-trained backbones on medical datasets. The top 3 results are highlighted with best in green and bold, light green, and third best in lightest green.
Figure 1: Variation of accuracy of top 3 backbones with increasing training data. The model performance at 3 orders of training data, 1% (500 images), 10% (5000 images) and 100% (50,000 images).
Figure 2: Variation of accuracy of top 3 backbones with increasing training data. The model performance at 3 orders of training data, 1% (1000 images), 10% (10,000 images) and 100% (100,000 images).
CIFAR-10 Tiny ImageNet
Number of training images 50,000 5,000 500 100,000 10,000 1,000
Percentage of training data 100% 10% 1% 100% 10% 1%
ResNet 94.58 86.5 65.42 76.30 66.56 28.11
WaveMix 97.26 93.16 80.98 75.33 55.15 25.24
ConvNeXt 96.48 91.06 72.22 83.42 67.47 37.38
Swin 95.38 85.64 39.77 76.85 48.50 25.62
SwinV2 94.52 85.66 46.51 73.93 39.31 12.29
EfficientNet 96.42 89.44 77.06 81.35 72.97 48.42
DenseNet 96.03 86.92 64.83 75.03 51.89 18.00
MobileNet 95.92 88.85 61.27 76.74 62.70 31.08
RegNet 96.82 90.55 73.71 80.14 64.40 36.56
ResNeXt 96.16 90.24 73.12 77.86 67.01 34.40
ShuffleNet 96.30 90.62 77.45 76.43 61.65 29.89
Table 7: The variation of accuracy of backbones with reduction of training data. Performance at 100%, 10% and 1% of training data for CIFAR-10 and Tiny ImageNet datasets
PlantVillage PlantCLEF RSNA Food-101 EuroSAT
Number of training images 44,343 4,434 10,455 1,045 24,181 2,418 75,750 7,575 18.900 1,890
Percentage of training data 100% 10% 100% 10% 100% 10% 100% 10% 100% 10%
ResNet 99.83 98.07 80.60 57.47 87.3 81.26 88.36 74.26 98.75 96.49
WaveMix 99.79 97.39 79.52 57.64 86.94 80.93 85.88 68.86 98.96 96.48
ConvNeXt 99.80 98.92 82.71 64.68 86.71 80.03 89.43 78.47 98.74 97.29
Swin 99.70 97.75 79.03 48.57 70.36 68.22 88.81 76.21 98.52 95.95
SwinV2 99.76 97.72 78.16 21.47 86.38 67.63 89.04 74.17 98.50 96.48
EfficientNet 99.81 98.42 81.09 64.4 87.06 79.96 89.84 76.68 98.88 97.23
DenseNet 99.88 98.57 76.49 58.59 87.1 81.51 87.02 72.44 98.83 96.86
MobileNet 99.80 98.29 75.52 54.93 87.05 79.55 85.73 71.12 98.72 96.35
RegNet 99.83 98.25 82.76 64.53 87.05 79.56 90.4 77.89 98.69 96.49
ResNeXt 99.75 98.01 80.32 63.61 87.89 81.42 88.89 74.69 98.96 96.53
ShuffleNet 99.70 98.30 81.76 58.69 87.53 80.42 86.21 70.97 98.65 96.00
Table 8: The variation of accuracy of backbones with reduction of training data. Performance at 100% and 10% training data.
@@ 7 Observations
$\bigtriangleup$ WaveMix performs well in datasets where multi-resolution token-mixing aid in learning WaveMix outperforms all other models (9% increase from the second best, ConvNeXt) in galaxy morphology classification. WaveMix also performs well in medical domain performing better than all other datasets, and also maintaining the performance across different magnification. WaveMix, which uses 2D-DWT might possess inductive bias that can analyse the domains of astronomy and medical images better than other convolutional models due its multi-resolution token-mixing. We recommend using WaveMx in domains where features across different resolutions are needed for better performance. WaveMix also performs better than ConvNeXt in CIFAR-10 and CIFAR-100 datasets, which were actually low resolution natural images ($32\times 32$) which were resized to ($256\times 256$) for our training. Similarly, it gives best performance in remote sensing dataset, EuroSAT, whose images ($64\times 64$) were resized to ($256\times 256$) for our training. WaveMix is also state-of-the-art in many low-resolution datasets such as EMNIST ($28\times 28$). The only low resolution image dataset where WaveMix did not perform well is Tiny ImageNet whose images ($64\times 64$) were also resized to ($256\times 256$). We attribute this to the fact that Tiny ImageNet is an ImageNet-1k subset and the other models which has better performance in ImageNet-1k naturally performed better. So, we recommend using WaveMix for low-resolution image datasets.
@@
$\bigtriangleup$ ConvNeXt and Swin transformers perform poorly in medical domain. Our results show that both the architectures perform perform really bad compared to all the other models when used in medical domain. Therefore, we suggest practitioners working in medical domain to not use these models despite their high performance in other domains.
$\bigtriangleup$ Most of the top models in every domain retain their higher performance even with less training data. We find that even when we fine-tune with a small percentage of training data (even 1% $\sim 1000$ images), the models which performed well with full training set still retained their superiority. This points to the presence of a domain specific inductive bias present in these models since they can learn better representations with very less data. The failure of these models to perform well on other domains similarly with less training data also alludes to this inductive bias.
Backbone Block
Kernels
in block
Stages
#Blocks/stage Non-linearity CEF Stem DConv SE Norm
ResNet-50 Bottleneck 1$\times$1, 3$\times$3, 1$\times$1 4 3,4,6,3 ReLU 4 7$\times$7, /2 No No BN
DenseNet-161 Inverted BN 1$\times$1, 3$\times$3 4 6,12,36,24 ReLU 4 7$\times$7, /2 No No BN
ResNeXt-50-32$\times$4d Bottleneck 1$\times$1, 3$\times$3, 1$\times$1 4 3,4,6,3 ReLU 2 7$\times$7, /2 No No BN
ShuffleNetV2-$\times$2.0 Isotropic 1$\times$1, 3$\times$3, 1$\times$1 3 4,8,4 ReLU 1 3$\times$3, /2 Yes No BN
MobileNetV3-Large Inverted BN 1$\times$1, 3$\times$3/5$\times$5, 1$\times$1 4 2,3,6,3 ReLU, SiLU 3-6 3$\times$3, /2 Yes Yes BN
RegNet-Y3-2GF Isotropic 1$\times$1, 3$\times$3, 1$\times$1 4 2, 5, 13, 1 ReLU 1 3$\times$3, /2 No Yes BN
EfficientNet-V2-S Inverted BN 1$\times$1, 3$\times$3, 1$\times$1 6 2, 4, 4, 6, 9, 15 SiLU 4-6 3$\times$3, /2 Yes Yes BN
WaveMix-192/16 Tokenmixer 1$\times$1, 1$\times$1, 1$\times$1 1 16 GELU 2 4$\times$4, /4 No No BN
ConvNeXt-Tiny Tokenmixer 7$\times$7, 1$\times$1, 1$\times$1 4 3,3,9,3 GELU 4 4$\times$4, /4 Yes No LN
Table 9: Architectural comparison of all the convolutional backbones showing the block structure of the residual block (eg. BottleNeck (BN)), kernels used in each block in sequence, number of stages in pyramidal models, number of blocks in each stage in pyramidal models, the non-linearities used in networks blocks, the channel expansion factor (CEF) in each block, the kernel and stride used in the initial stem layer, whether depth-wise convolutions (DConv) and squeeze-excitation (SE) operations were used in blocks, and the normalization used in blocks (whether BatchNorm (BN) or LayerNorm (LN)). The architectures are ordered in descending order of age.
8 Architectural Discussions
Table 9 shows the architectural details of all the convolutional architectures used in our experiments from oldest to latest. From the Table, we can see certain architectural trends that has improved the performance of the newer models compared to older ones. All top performing models in all the domains are the latest models such as RegNet, EfficientNet, WaveMix and ConvNeXt. These observations can be used to develop better resource-efficient backbones for fine-tuning tasks.
Most of the models use a $1\times 1$, $3\times 3$, $1\times 1$ kernel structure in their blocks with a normalisation and activation following each convolutional operation. MobileNet uses $5\times 5$ kernels towards the final stages, while ConvNeXt uses $7\times 7$ kernel for spatial token-mixing in each block. All architectures use $1\times 1$ convolution to change the number of channels.
While increasing the channel dimensions inside bottlenecks/inverted bottlenecks, almost all models use a channel expansion/contraction factor of 4. MobileNet and EfficientNet uses even higher values up to 6.
All models use adaptive average pooling in the classification head to down-sample the output feature maps from the final stage to $1\times 1$ spatial resolution.
$\bigtriangleup$ Inverted bottleneck design is better. The older models used bottleneck blocks where channel dimension was reduced using $1\times 1$ convolutions before being passed to the main $3\times 3$ convolutions to increase efficiency. Newer architectures such as EfficientNet and MobileNet use an inverted bottleneck structure where the $1\times 1$ convolutions increases the channel dimension before $3\times 3$ convolutions. The increase in parameters is offset mainly by using the parameter efficient depth-wise convolution which also increases performance. Models like ShuffleNet and RegNet keeps the channel dimension invariant throughout the block resulting in an isotropic design. WaveMix and ConvNeXt follow the token-mixing structure where spatial and channel mixing is performed separately, with spacial mixing done with token-mixing operations such as wavelet transform and large kernel depth-wise convolution respectively. The channel dimension is expanded only for channel mixing using $1\times 1$ convolutions in an inverted bottleneck design.
$\bigtriangleup$ Depth-wise convolution replacing regular convolution. Depth-wise convolutions have fewer parameters and requires less operations than regular convolutions. Most of the newer models such as ShuffleNet, MobileNet and EfficientNet replace the $3\times 3$ kernel regular convolution with $3\times 3$ kernel depth-wise convolution. ConvNeXt uses $7\times 7$ kernel depth-wise convolution for efficient spacial token-mixing. All $1\times 1$ convolutions in all these models are point-wise convolutions (regular convolutions). SiLU activation has been observed to give better performance when used with depth-wise convolutions [17].
$\bigtriangleup$ 4 stage models with largest number of blocks in penultimate stage. From old to new models, 4 stage structure has been most popular and has resulted in good performance. Other than EfficientNet with 6 stages, ShuffleNet with 3 stages and WaveMix with a single stage, all other backbone has 4 stages. The design principle of increasing the number of blocks with each subsequent stage with the exception of final stage is also observed in most models. Only Efficient has more blocks in the final stage than the penultimate stage. Most of the 4 stage models use a channel expansion factor greater than or equal to 2 during feature resolution down-sampling between stages to increase the number of channels.
$\bigtriangleup$ Strided convolution for down-sampling. With the exception of DenseNet which used average pooling for down-sampling the feature resolution between different stages of the network, all other convolutional models used strided convolutions with stride 2 for feature resolution down-sampling. So strided convolutions are better for model performance than pooling operations.
$\bigtriangleup$ Small kernel stem layer. Stem layer is the initial layer used to down-sample the full size input image whose output is then sent to the body of architectures which has multiple stages. Compared to older models which used large $7\times 7$ kernel with stride 2, newer models use $3\times 3$ and $4\times 4$ kernels. The token-mixers use patchify stem ($4\times 4$ non-overlapping convolutions).
$\bigtriangleup$ SiLU and GELU replacing ReLU. Older models used ReLU activation after convolution operations. The newer models use a combination of ReLU and SiLU activations while the token-mixers inspired from transformer architecture uses GELU activation. Compared to other convolutional models which uses 2 to 3 activations per block, token-mixers use much less activations (one per block).
$\bigtriangleup$ Addition of squeeze-excitation (SE) module. [39] SE module enhance CNNs by focusing on the most informative features through explicit modeling of channel interdependencies and can be easily integrated into various architectures without significant changes. The top performing architectures such as RegNet and EfficientNet used SE modules to improve performance.